You have two SNPs each with two alleles that have been implicated in Parkinson’s Disease.

AG=687 AC=805 TG=209 TC=943

A) Calculate the haplotype and allele frequencies

B) Calculate the D, D’ and r2 for the SNPs

C) Are the SNPs in significant LD? Use Chi Square to determine.
Answer #1

Given catas, AG=687 AC =805 C =805 TG = 209 TG = 209 TC = 943 (A) Total (N)= AGT Act TGATC = 687 +805+ 209 +943 = 2644 Haplot

s2 for the SNPsi- r2 = D/Spix p2x91x991/2)) q=1.99964% POC g 2 = 3.9992 (C) xax N = 84.0792